OpenStudy (anonymous):

How could i solve this? please help me! There are 20 coins altogether. and pennies and nickels were the only coins used. The total value is 48 cents. how do i find out how many pennies and nickels were used?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (austinl):

This will be a system of equations problem. If that helps.

OpenStudy (anonymous):

well i know there is 7 nickels and 13 pennies, I found this out by using a different method but i need to show it using two different methods, they were talking about other methods like system of linear equations and solve by graphing

OpenStudy (anonymous):

which one out of those two is easier to solve?

OpenStudy (anonymous):

or do you know of any easier methods?

OpenStudy (austinl):

\(P+N=20\\ P=\text{Pennies}\\ N=\text{Nickles}\) \(0.01P+0.05N=0.48\) So with these we have the system: \(P+N=20\) \(0.01P+0.05N=0.48\)